# Symplectic Lattices and GKP Codes - Simple Randomized Constructions from Cryptographic Lattices

### v1

@@ 10 Main results - module and ring constructions
In this section we use Theorem 32 and Corollary 33 in combination with Theorem 23 to construct $n$-dimensional symplectic lattices with minimal distance $\approx\sqrt{n/\pi e}$, and then, as in Section 5, good GKP codes. Theorem  32 and Corollary 33 only yield meaningful results, if the right-hand sides in equations ( 21) and ( 23) do not contain to many terms. Equivalently, the factorization in Theorem 23 may not contain too many terms. When $X^{n}+1$ factors into a product of linear terms, this is certainly not the case [^4] Note that this case is interesting in many cryptographic applications of R-SIS lattices. Since then the number of terms in (21) in exponential in $n$. Fortunately, Theorem 23, together with (13), offers us interesting and useful alternatives. More precisely, the results we present below concentrate on choices of $n,q,r$ such that
  1. 1.
    the number of terms in the sums of the right-hand sides of (21) and  (23) are small,
  2. 2.
    for all subsets $\cal T$}, except possibly one, the degrees $d_{\cal T}$ are large, i. e. $d_{\cal T}=\Omega(n^{\delta}),\delta>$, and for the exceptional set $\cal T$} the degree $d_{\cal T}$ is a power of $2$.
We consider the case $n=2^{e}p_{1}^{e_{1}}p_{2}^{e_{2}},$ where $p_{1},p_{2}$ are distinct odd primes. Then, according to Theorem 12, the factor of $X^{n}+1$ in $\mathbb{Z}$ are $\Phi_{2d}(X),$ where
$2d\in\{2^{e+1}p_{1}^{d_{1}}p_{2}^{d_{2}}\mid 0\leq d_{1}\leq e_{1},0\leq d_{2}\leq e_{2}\}.$
To apply Theorem 23 with some primitive $\lambda$-root $q$ of $2n$, and hence primitive $\lambda$-root for all $2d,d|n$, we need determine the values $\phi(2d),\lambda(2d),$ and $\phi(2d)/\lambda(2d),$. To do so, for $2d=2^{e+1}p_{1}^{d_{1}}p_{2}^{d_{2}}$, write
$\phi(p_{i}^{d_{i}})=p_{i}^{d_{i}-1}(p_{i}-1))=2^{f_{i}}u_{i}p_{i}^{d_{i}-1},\;\text{$f_{i}\geq 1,u_{i}$ odd},$
and set $u=\gcd(p_{1}^{d_{1}-1}u_{1},p_{2}^{d_{2}-1}u_{2})$. Then, by straightforward calculations,
$\lambda(d)=\begin{cases}2^{\max\{e,f_{1},f_{2}\}}\frac{\phi(p_{1}^{d_{1}}\phi(p_{2}^{d_{2}})}{2^{f_{1}+f_{2}}u}&e=0,1\\ 2^{\max\{e-1,f_{1},f_{2}\}}\frac{\phi(p_{1}^{d_{1}}\phi(p_{2}^{d_{2}})}{2^{f_{1}+f_{2}}u}&e\geq 2.\end{cases}$ (24)
Hence
$\frac{\phi(2d)}{\lambda(2d)}=\begin{cases}2^{e+f_{1}+f_{2}-\max\{e,f_{1},f_{2}\}}\cdot u&e=0,1\\ 2^{e+f_{1}+f_{2}-\max\{e-1,f_{1},f_{2}\}}\cdot u&e\geq 2.\end{cases}$ (25)
These formulas also apply if $d_{1}=0$ or $d_{2}=0$. For the two conditions form above to hold we want
  1. 1.
    $e_{1},e_{2}$ to be small (this limits the number of factors $\Phi_{2d}(X)$ of $X^{n}+1$ over $\mathbb{Z}$)
  2. 2.
    for all $2d$, the values $f_{1},f_{2},u$ defined above to be small (this limits the number of factors of $\Phi_{2d}(X)$ modulo $q$).
The first, and most important, of our constructions construction uses $n=2^{e}$. This setting is also by Conrad et al. in  [18]. But whereas Conrad et al. use NTRU lattices we use M-SIS lattices from cryptography. Other possible choices for $n,q$ are discussed below. To choose appropriate values for $r$ and analyze the upper bounds for the probability of small minimal distance in Theorem 32 and Corollary 33, we concentrate on subsets $\cal T$} leading to polynomials with smallest degree $d:=d_{\cal T}$ and consider the term
$\operatorname{smd}(n,k,q,r,d):=\frac{1}{\sqrt{2\pi kd}}\left(\frac{\pi e}{nk}\right)^{kd}\left(r+\sqrt{\frac{nk}{2q}}\right)^{2kd}.$ (26)
Setting
$\rho=\rho(n,k,\gamma):=r+\sqrt{\frac{nk}{2q}}=e^{-\gamma/2}\sqrt{\frac{nk}{\pi e}},\;\text{for some}\;\gamma\in\mathbb{R},$ (27)
we get
$\operatorname{smd}(n,k,q,r,d)=\frac{1}{\sqrt{2\pi dk}}e^{-\gamma dk}\;\text{and}\;r\geq\sqrt{\frac{nk}{\pi e}}\left(e^{-\gamma/2}-\sqrt{\frac{\pi e}{2q}}\right).$ (28)
Finally, set
$q\geq\frac{2\pi e}{\gamma^{2}}.$ (29)
Using, $e^{x}>1+x$ for all $x$, we get
$r>\begin{cases}\sqrt{\frac{nk}{\pi e}}(1-\gamma),&\gamma\geq 0\\ \sqrt{\frac{nk}{\pi e}},&\gamma<0.\end{cases}$ (30)
With these settings and bounds first we get the following results.
@@
    For $q\geq\frac{\pi e(nk)^{2}}{2}$, with probability at least $1-\frac{e^{2}+2\sqrt{2}e}{\sqrt{2\pi nk}}=1-\Omega(1/\sqrt{nk})$
@@
    For $q\geq 2\pi e(nk)^{3/2}$, with probability at least $1-e^{-(nk)^{1/4}}$
@@
The proof of Theorem 2’ in Chapter 4 of [32] shows that for $q=5\mod 8,n\geq 8$, the order of $q$ modulo $2n$ is $n/2$. That is, $q$ is a primitive $\lambda$-root modulo $n$. The same proof can be used for $q=3\mod n$ and $n\geq 16$.
By Theorem 23, Equation (25) and with our choice of $q$, modulo $q$ polynomial $X^{n}+1$ factors into two polynomials $p_{1}(X),p_{2}(X)$ of degree $n/2$ each. Hence the right-hand side in (23) has three summands for the sets ${\cal T}=\{1\},\{2\},\{1,2\}$. The corresponding degrees $d_{\cal T}$ are given by $n/2,n/2,n$ with values $g_{\cal T}$ and $n/g_{\cal T}$ given by $n/2,n/2,n$ and $2,2,1,$ respectively. Using Corollary 33 and the notation and bounds from above, we obtain
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ 2\cdot\operatorname{smd}(n,k,q,r,n/2)+\operatorname{smd}(n,k,q,r,n).$ (31)
  1. 1.
    In this case, using the notation from above, we set $\gamma:=-2/kn$ and, correspondingly, choose a prime $q\geq\pi e(kn)^{2}/2$ (see (29)). This yields
@@
    and
[…]
  2. 2.
    Here we choose
    $\gamma:=\frac{2}{(nk)^{3/4}}\;\text{and}\;q\geq\frac{\pi e(nk)^{3/2}}{2},$
    and use the same line as arguments as in the previous case.
Next we consider the case where $n$ is the product of $2$ and two odd primes.
@@
[…]
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H}).$
  1. 1.
    For $k>\ln(n/2)/4$ and $q\geq\frac{32\pi ek^{2}}{\ln(n/2)^{2}}$, with probability at least $1-\frac{1+e^{-\Omega(p_{1})}}{\sqrt{2\pi nk}}=1-\Omega(1/\sqrt{nk})$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{\ln(n/2)}{4k}\right)\sqrt{\frac{nk}{\pi e}}.$
  2. 2.
    For $4k>n^{1/4}$ and $q\geq\frac{32\pi ek^{2}}{\sqrt{n}}$, with probability at least $1-(e^{-p_{1}^{1/4}}+e^{-\Omega(p_{1})}$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{n^{1/4}}{4k}\right)\sqrt{\frac{nk}{\pi e}}.$
Proof.
From Theorem 23 and Equations (24), (25) we get that modulo $q$ the polynomial $X^{n}+1$ factors into the polynomial $X^{2}+1$, two polynomials of degree $p_{1}-1$ (factors of $\Phi_{4p_{1}}(X)$ modulo $q$), two polynomials of degree $p_{2}-1$ (factors of $\Phi_{4p_{1}}(X)$ modulo $q$), and four polynomials of degree $\frac{(p_{1}-1)(p_{2}-1)}{2}$ (factors of $\Phi_{4p_{1}p_{2}}(X)$ modulo $q$). Hence, the right-hand sum in Equation 23 consists of $511$ terms. The smallest degree $d_{\cal T}$ is $2$ and the next smallest degree is $p_{1}-1$. The (asymptoktically) determining term corresponds to the polynomials $X^{2}+1$. In which case we can use Remark 34 and focus on $\operatorname{smd}(n,k,q,r,2)$.
  1. 1.
    We choose
    $\gamma:=\frac{\ln(n/2)}{4k}\quad\text{and}\quad q\geq\frac{\pi e32k^{2}}{\ln(n/2)}.$
    As in the proof of the previous theorem, we can show that with the choices
    $\operatorname{smd}(n,k,q,r,2)\leq\frac{1}{2\pi nk}\quad\text{and}\quad r\geq\left(1-\frac{\ln(n/2)}{4k}\right)\sqrt{\frac{nk}{\pi e}}.$
    For all other sets $\cal T$} with corresponding degree $d=d_{\cal T}$ we have $d\geq p_{1}-1$ and hence
    $\operatorname{smd}(n,k,q,r,d)<e^{-\gamma kd}\leq e^{-p_{1}/8}.$
    This implies that for the other set $\cal T$} their combined contribution to $\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]$ is bounded by
    $511n\operatorname{smd}(n,k,q,r,d)=e^{-\Omega(p_{1})}.$
    This proves the first part of the theorem.
  2. 2.
    The proof of the second part of the theorem follows as the proof for the first part by choosing
    $\gamma=\frac{n^{1/4}}{4k}\quad\text{and}\quad q\geq\frac{32\pi ek^{2}}{n^{1/2}}.$
Finally, consider the case $n=2^{e}p,e\geq 2,p$ prime. We only state the result. Its proof follows along the same lines as the previous two proofs.
@@
Let $k,n,q\in\mathbb{N}$ with $n=2^{e}p,e\geq 2$, and $p,q$ primes such that $q$ is a primitive $\lambda$-root modulo $2n$. Set $R:=\mathbb{Z}[X]/(X^{n}+1)$ and $R_{q}:=\mathbb{Z}[q]/(X^{n}+1)$. Choose H uniformly at random from $R_{q,k}^{\operatorname{Sym}}$ and set
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H}).$
  1. 1.
    For $k>\ln(n/2^{e})/2^{e+1}$ and $q\geq\frac{2^{2e+3}\pi ek^{2}}{\ln(n/2^{e})^{2}}$, with probability at least $1-\frac{1+e^{-\Omega(n)}}{\sqrt{2\pi nk}}=1-\Omega(1/\sqrt{nk})$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{\ln(n/2^{e})}{2^{e+1}k}\right)\sqrt{\frac{nk}{\pi e}}.$
  2. 2.
    For $2^{e+1}k>n^{1/4}$ and $q\geq\frac{2^{2e+3}\pi ek^{2}}{\sqrt{n}}$, with probability at least $1-(e^{-n^{1/4}}+e^{-\Omega(n)})$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{n^{1/4}}{2^{e+1}k}\right)\sqrt{\frac{nk}{\pi e}}.$
[…]
@@ 11 Bounded distance decoding in the module case and the role of $k$
We can also use algorithm $\textsc{Bdd}_{\text{triv}}$ from Section 6 to decode the GKP codes obtained from M-SIS lattices. The most time consuming step in algorithm $\textsc{Bdd}_{\text{triv}}$ is the matrix-vector multiplication c’H in step 2, which requires $n^{2}$ multiplications in $\mathbb{F}_{q}$. If $n$ is a power of $2$, instead of the SIS lattice construction of GKP codes we can use the M-SIS-based constructed using Theorem 35, with $k=1$. The matrix-vector multiplication in step $2$ of algorithm $\textsc{Bdd}_{\text{triv}}$ becomes a multiplication of two ring elements in $R=Z[X](X^{n}+1)$. By construction, this ring contains $n$-th roots of unity which allows us to perform multiplication in $R$ with ${\textbf{{O}}}(n\log(n))$ arithmetic operations in $\mathbb{F}_{q}$ using the fast Fourier transform (FFT).
However, powers of $2$ are relatively sparse. Using $k>1$, allows us to obtain fast decoding for more general $n$. If we choose the dimension of the code as $n\cdot k$ with $n$ being a power of $2$, the multiplication c’H requires $k^{2}$ multiplications in ring $R=\mathbb{Z}[X]/(X^{n}+1)$ at the cost ${\textbf{{O}}}(k^{2}n\log(n))$ arithmetic operations in $F_{q}$, again using FFT. Hence, the parameter $k$ and the use of M-SIS lattices and not just R-SIS lattice (corresponding to $k=1$) allows us more flexibility in designing code GKP codes with efficient decoding algorithms. Using other choices for $n$ than just powers of $2$, yields yet more flexibility in designing good GKP codes for various dimensions.

### v2

@@ 10 Main results - module and ring constructions
In this section, based on the analysis of the previous section, we prove lower bounds on the minimum distance of random $q$-ary lattices, focussing on two types of bounds: a lower bound of $\sqrt{nk/\pi e}$ exactly and lower bounds close to $\sqrt{nk/\pi e}$, that can be obtained with probability exponentially close to $1$. We begin with the strongest and most interesting case, namely $n=2^{e}$ being a power of $2$.
@@
    For $q\geq 11\pi e(nk)^{2}$, with probability at least $1-\frac{6}{\sqrt{nk}}$
@@
    For $q\geq 2\pi e(nk)^{3/2}$, with probability at least $1-e^{-(nk)^{1/4}/2}$
@@
Note that $\lambda_{1}({\cal L})=\frac{1}{\sqrt{q}}\lambda_{1}({\cal L}^{\perp}(\textbf{H}))$. Hence for both statements it suffices the prove the statements for ${\cal L}^{\perp}(\textbf{H})$ with bounds $\sqrt{q}\lambda_{1}({\cal L})$.
To prove the first statement, set $p:=6/\sqrt{nk}$. Then
$\left(\frac{p\sqrt{kn}}{4}\right)^{\frac{1}{nk}}=(3/2)^{\frac{1}{nk}}\geq 1+\frac{\ln(3/2)}{nk}.$
With our choice of $q$ it follows that
$\sqrt{\frac{nk}{2q}}\leq\frac{\ln(3/2)}{nk}\sqrt{\frac{nk}{\pi e}}.$
The statement follows from Theorem 33.
To prove the second statement we proceed similarly. First, with our choice of $p$ we obtain
$\left(\frac{p\sqrt{kn}}{4}\right)^{\frac{1}{nk}}\geq e^{(nk)^{3/4}/2}\geq\left(1-\frac{1}{2(nk)^{3/4}}\right).$
By choice of $q$
$\sqrt{\frac{nk}{2q}}\leq\frac{1}{2(nk)^{3/4}}.$
As before, the statement follows from Theorem 33. ∎
For the general case of $n,q$ and the factorization of $X^{n}+1$ modulo $q$, a lower bound on the minimum distance of a random $q$-ary lattice can achieved with high probability if $d$ is large enough.
Theorem 35.
Let $k,n,q\in\mathbb{N}$ with $q$ prime and let $d$ be the minimum degree of a polynomial in the factorization of $X^{n}+1$ modulo $q$. Let $0<p,\gamma<1$. If
$d\geq\frac{\ln(1/p)+2\ln(n)}{2\gamma k}\qquad\text{and}\qquad q\geq\frac{2\pi e}{\gamma^{2}}$
then for H uniformly at random from $R_{q,k}^{\operatorname{Sym}}$ and ${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ with probability at least 1p1-p
$\lambda_{1}({\cal L})\geq(1-\gamma)\sqrt{\frac{nk}{\pi e}}.$
Proof.
We proceed as in the proof of the previous theorem, using Theorem 31 instead of Theorem 34. From the lower bound on $d$, it follows that
$\left(\frac{p\sqrt{kd}}{n\sum_{m}n_{m}}\right)^{\frac{1}{2kd}}\geq e^{-\gamma/2}\geq(1-\gamma/2).$
The bound on $q$ implies
$\frac{nk}{2q}\leq\gamma/2\frac{nk}{\pi e}.$
The theorem follows from Theorem 31. ∎
It is not difficult to show that, using Theorem 31, the bound in the previous theorem cannot be improved significantly.
Instantiations
First, we show how to calculate $d:=\min\{d_{m}\}$ given specific values of $n$ and $q$. To do so, we need to recall some facts from elementary number theory, that yield more details on the factorization of $X^{n}+1$ modulo $q$, (see (12)).
We use the Carmichael function $\lambda(\cdot)$, where for $n\in\mathbb{N}$, $\lambda(n)$ is the smallest integer $m$ such that
$\forall a\in\mathbb{Z}_{n}^{*}:a^{m}=1\mod n.$
@@
Lemma 36.
@@
This theorem implies for a primitive $\lambda$-root $q$ of $n$:
$d:=\min\{d_{m}\}:=\min\{\lambda(m):m\mid 2n\land m\nmid n\}$
Write $n=2^{e}n^{\prime},$ with $n^{\prime}$ odd. Then this characterization of $d$ together with (28) implies
$d=\begin{cases}2^{e}&e=0,1\\ 2^{e-1}&e\geq 2.\end{cases}$
If $n=2^{e}$ is a power of $2$ and $q=5\mod 8$ or $q=3\mod 16$, then $q$ is a primitive $\lambda$-root modulo $q$. Hence, as already used in Theorem 34, $d=n/2$.
Theorem 35 can give meaningful results even for $d=1$. In particular, if we exploit the general module structure, i.e. we may choose $k$ freely, although we want to keep it as small as possible. As an example, in Theorem 35, let us choose $d=1$ and $p=1/n$. Then we get
$k\geq 3\ln(n)/(2\gamma).$
Hence by increasing $k$ we can get, with probability close to $1$, GKP codes whose minimum distance gets closer to $\sqrt{nk/\pi e}$.
If on the other hand, we want to choose $k=1$, i.e. we choose ring lattices, not general module lattice, then with $p=1/n$ Theorem 35 implies
$d\geq\frac{3\ln(n)}{2\gamma},$
hence $d$ must be a least logarithmic.
One example where $d=1$, in fact, where all factors of $X^{n}+1$ modulo $q$ are linear, occurs when $q=1\mod n$. This case is quite popular in lattice-based cryptography, since it allows for fast arithmetic in $R_{q}$ using the so called number theoretic transformation (NTT), a variant of the Fourier transform. As we will see in the next section, the case $q=1\mod n$ also can lead to efficient decoding algorithms for GKP codes.
11 GKP codes from NTRU lattices
In this section we combine the results from Section 8 with results from [32] to show that NTRU lattices can be used to construct good GKP codes, thereby proving a conjecture by Conrad et al. [11].
For the exact definition of the NTRU cryptosystem we refer to [19]. By $R_{q,\sigma}^{\text{NTRU}}$ we refer to the following distribution, which we call the $q$-ary NTRU distribution with parameter $\sigma$ or simply the NTRU distribution. First, denote by $D_{\sigma}^{\times}$ the discrete Gaussian distribution $D_{\mathbb{Z}^{n},\sigma}$ restricted to $R_{q}^{\times}$ (via rejection sampling). Then $R_{q,\sigma}^{\text{NTRU}}:=\frac{D_{1}}{D_{2}}\mod q,$ where $D_{1},D_{2}$ are independently distributed as the Gaussian distribution $D^{\times}_{\sigma}$ Note that $R_{q,\sigma}^{\text{NTRU}}$ is a distribution on $R_{q}^{\times}$. By $U(R_{q}^{\times})$ and $U(R_{q})$ we denote the uniform distributions on $R_{q}^{\times}$ and $R_{q}$, respectively.
For the remainder of this section let $n=2^{e}$ be a power of two and $q=5\mod 8$ or $q=3\mod 16$. In this case $X^{n}+1$ is irreducible over $\mathbb{Q}$ and it factors into two polynomials of degree $n/2$ over $\mathbb{F}_{q}$. From this it follows using the Chinese remainder theorem that
$\left|R_{q}^{\times}\right|=(q^{n/2}-1)^{2}\qquad\text{and}\qquad\left|R_{q}\setminus R_{q}^{\times}\right|<2q^{n/2}.$ (29)
For two distributions $D_{1},D_{2}$ with identical finite support $T$, denote by $\Delta(D_{1},D_{2})$ their statistical distance, that is
$\Delta(D_{1},D_{2}):=\max\{|D_{1}(S)-D_{2}(S)|\mid S\subseteq T\}=\frac{1}{2}\sum_{t\in T}|D_{1}(t)-D_{2}(t)|.$
From (29) one deduces
$\Delta(U(R_{q}),U(R_{q}^{\times}))\leq\frac{2q^{n/2}}{q^{n}-2q^{n/2}}\leq\frac{3}{q^{n/2}},\qquad\text{$q>5$}.$ (30)
As a special case of Theorem 3.2 in [32], we obtain the following theorem.
@@ 11 GKP codes from NTRU lattices
[Stehlé, Steinfeld] Let $n=2^{e}$ be a power of two and $q=5\mod 8$ or $q=3\mod 16$. Let $0<\epsilon<1/4$. Then
$\Delta\left(R_{q,\sigma}^{\text{NTRU}},U(R_{q}^{\times})\right)\leq 2^{10n}q^{-\epsilon n},\;\text{if}\;\sigma\geq\sqrt{n\ln(8n)}q^{1/2+\epsilon}\;\text{and}\;q\geq n^{\frac{2}{1-4\epsilon}}.$
Combining this theorem with (30) yields
Corollary 39.
Let $n=2^{e},n>5,$ be a power of two and $q=5\mod 8$ or $q=3\mod 16$. Let $0<\epsilon<1/3$. Then
$\Delta\left(R_{q,\sigma}^{\text{NTRU}},U(R_{q})\right)\leq(2^{10n}+3)q^{-\epsilon n},\;\text{if}\;\sigma\geq\sqrt{n\ln(8n)}q^{1/2+\epsilon}\;\text{and}\;q\geq n^{\frac{2}{1-4\epsilon}}.$
For $k=1$, we have that $R_{q,k}^{\operatorname{Sym}}=U(R_{q})$. Hence, combining Corollary 39 with part (2) of Theorem 34 leads to the following theorem, proving Conjecture 2 from [11], at least asymptotically.
Theorem 40.
Let $n=2^{e},n>5,$ be a power of two and $q=5\mod 8$ or $q=3\mod 16$. Let $0<\epsilon<1/4$. Choose h from the NTRU distribution $R_{q,\sigma}^{\text{NTRU}}$, where $\sigma\geq\sqrt{n\ln(8n)}q^{1/2+\epsilon}$ and $q\geq\max\{11\pi 2n^{3/2},n^{\frac{2}{1-4\epsilon}}\}$. Then
$\lambda_{1}({\cal L})\geq\left(1-\frac{1}{(n)^{3/4}}\right)\sqrt{\frac{n}{\pi e}},$
except with probability $e^{-n^{1/4}/2}+(2^{10n}+3)q^{-\epsilon n}$.
@@ 12 Bounded distance decoding in the module case and the role of $k$
We can also use algorithm $\textsc{Bdd}_{\text{triv}}$ from Section 6 to decode the GKP codes obtained from M-SIS lattices. The most time consuming step in algorithm $\textsc{Bdd}_{\text{triv}}$ is the matrix-vector multiplication c’H in step 2, which requires $n^{2}$ multiplications in $\mathbb{F}_{q}$. If $n$ is a power of $2$, instead of the SIS lattice construction of GKP codes we can use the M-SIS-based construction using Theorem 34, with $k=1$. The matrix-vector multiplication in step $2$ of algorithm $\textsc{Bdd}_{\text{triv}}$ becomes a multiplication of two ring elements in $R=Z[X](X^{n}+1)$. By construction, this ring contains $n$-th roots of unity which allows us to perform multiplication in $R$ with ${\textbf{{O}}}(n\log(n))$ arithmetic operations in $\mathbb{F}_{q}$ using the fast Fourier transform (FFT).
However, powers of $2$ are relatively sparse. Using $k>1$ allows us to obtain fast decoding for more general $n$. If we choose the dimension of the code as $n\cdot k$ with $n$ being a power of $2$, the multiplication c’H requires $k^{2}$ multiplications in ring $R=\mathbb{Z}[X]/(X^{n}+1)$ at the cost ${\textbf{{O}}}(k^{2}n\log(n))$ arithmetic operations in $F_{q}$, again using FFT. Hence, the parameter $k$ and the use of M-SIS lattices and not just R-SIS lattice (corresponding to $k=1$) allows us more flexibility in designing code GKP codes with efficient decoding algorithms. Using other choices for $n$ than just powers of $2$, yields yet more flexibility in designing good GKP codes for various dimensions.